Write an equation of the line. Show all work.

slope = one third.

y-intercept= -4

thank you

Found 2 solutions by Alan3354, rapaljer:
Answer by Alan3354(69443)

There's no work to show.
y = (1/3)x - 4

Answer by rapaljer(4671)

There is a formula for the equation of a line that goes like this: y=mx+b, where m is the slope and b is the y intercept. So, if they give you the slope and y intercept, just plug the numbers into this equation:

y=mx + b
y= 1/3 x -4

For additional explanation on Graphing and Slope of a Line, please see my own website. Do a "Bing" search for my last name "Rapalje". Look for "Rapalje Homepage" near the top of the search list. On my Homepage near the top, look for the link "Basic, Intermediate and College Algebra: One Step at a Time." Choose "Basic Algebra", and look in Chapter 4 for several topics on GRAPHING. In these sections, you will find my own non-traditional explanations, that my own students all said were a LOT easier to understand than traditional textbooks. See also the corresponding "MATH IN LIVING COLOR" pages in which problems are solved IN COLOR! I also have FREE videos on the website of me teaching my classes a few years ago. To access the videos, go back to my Homepage, and look for the link "Rapalje Videos in Living Color." Select "Basic Algebra", and select the video on "GRAPHING and SLOPE of a LINE."

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus